A cyclist traveled at a rate of 32 mph to visit a nearby town. The cyclist averaged 10 mph on the return trip. If the round trip took 6.3 hours, find the distance to the nearby town.
Answer by ptaylor(2198)

Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
Let d=distance to nearby town
Time to nearby town=d/r=d/32
Time for the return trip=d/r=d/10
Now we are told that the above two times add up to 6.3 hrs, so:
d/32 + d/10=6.3 multiply each term by 160
5d + 16d=1008
21d=1008
d=48 mi---------------distance to nearby town
CK
time to nearby town=48/32=3/2 hrs
time to return =48/10 hrs
3/2+48/10=6.3
1.5+4.8=6.3
6.3=6.3
Hope this helps---ptaylor